### v2

# Portable, Massively Parallel Implementation of a Material Point Method for Compressible Flows

Abstract
The recent evolution of software and hardware technologies is leading to a renewed computational interest in Particle-In-Cell (PIC) methods such as the Material Point Method (MPM). Indeed, provided some critical aspects are properly handled, PIC methods can be cast in formulations suitable to the requirements of data locality and fine-grained parallelism of modern hardware accelerators as Graphics Processing Units (GPUs). Such a rapid and continuous technological development increases also the importance of generic and portable implementations. While continuum mechanics simulations have already shown the capabilities of MPM on a wide range of phenomena, the use of the method in compressible fluid dynamics is less frequent. In this paper we present a portable, highly parallel, GPU based MPM solver for compressible gas dynamics. The implementation aims to reach a good compromise between portability and efficiency and to give a first assessment of the potential of this approach in reproducing strongly compressible gas flows, also taking into account solid obstacles. The proposed model constitutes a first step towards the development of a monolithic MPM solver for Fluid-Structure Interaction (FSI) problems at all Mach numbers up to the supersonic regime.
(1) MOX – Modelling and Scientific Computing
Dipartimento di Matematica, Politecnico di Milano
Piazza Leonardo da Vinci 32, 20133 Milano, Italy
{paolojoseph.baioni,carlo.defalco}@polimi.it
(2) Danish Meteorological Institute
Sankt Kjelds Plads 11
2100 Copenhagen, Denmark
tbo@dmi.dk
(3) Leonardo Labs, Leonardo S.p.A.
Torre Fiumara, Via Raffaele Pieragostini, 80, 16149 Genova, Italy
luigi.capone@leonardo.com
Keywords: Material Point Method; GPUs; Compressible Gas Dynamics; Performance Portability; High-Performance Computing.
This work has not yet been peer-reviewed and is provided by the contributing authors as a means to ensure timely dissemination of scholarly and technical work on a noncommercial basis. Copyright and all rights therein are maintained by the authors or by other copyright owners. It is understood that all persons copying this information will adhere to the terms and constraints invoked by each author’s copyright. This work may not be reposted without explicit permission of the copyright owner. This work has been submitted for publication. Copyright in this work may be transferred without further notice.
1 Introduction and Motivation
Hardware and software technology evolution has led to a renewed interest in particle-based numerical techniques for differential problems, such as the Material Point Method (MPM, [51, 63]), that was developed stemming from the Particle In Cell method Fluid Implicit Particle Method [29, 8] to treat history-dependent materials, thanks to their suitability to modern hardware accelerators requirements [28, 22]. Indeed, parallel computing is a key capability for effective advancement of industrial-grade, large-scale computations, particularly in the fields of high-speed, compressible computational fluid dynamics and Fluid-Structure Interaction (FSI, [37]) simulations. On the other hand, rapid hardware and software evolution makes researchers lean towards generic and portable implementations, which enables developers to adapt fast to sudden technology innovations.
In view of exponentially increasing High Performance Computing (HPC) resources and the method’s accurate handling of a wide range of phenomena, MPM is becoming increasingly relevant for compressible CFD and FSI simulations. In MPM, continuum media are discretised by means of Lagrangian particles, the material points, each one storing every state variable evaluated at its position, enabling the description of a wide range of materials [26, 49, 32] as well as large deformations [1, 58, 18] and fracture phenomena [60]. Notwithstanding its Lagrangian character, MPM also uses a background Cartesian grid to compute differential quantities and solve the motion equation, thus mediating the particle-particle interactions, and taking advantage of both Eulerian and Lagrangian approaches. These features make it an interesting candidate for the development of a highly parallel, Graphics Processing Units (GPUs)-based FSI solver. Indeed, in order to reach a high degree of efficiency, GPU-based solvers require a large degree of data locality and fine-level parallelism. MPM implementations can naturally achieve these, since:
  • every GPU thread can manage up to very few grid cells or particles;
  • the link between cells and particles can be easily built thanks to the structured character of the grid;
  • accuracy in the discretisation of continuum bodies is obtained by means of the Lagrangian particles, so that the grid does not have to match the solid boundary.
Thus, there is no need of computationally expensive, unstructured, moving or deforming meshes and a structured Cartesian grid can be used instead, simplifying also techniques such as Adaptive Mesh Refinement and Domain Decomposition in a parallel context [53, 61].
Since MPM first introduction, various enhancements of the method were proposed, solving some of its drawbacks such as grid-crossing error [55] and volumetric locking [17], and improving its accuracy and conservation properties. To address instabilities which arise when particles cross the grid, variants of the method were proposed, such as GIMP [4], CPDI [45], iMPM [52] and IGA-MPM [40], and numerical analysis studies on the use of B-Splines as basis functions in MPM were conducted [48]. An affine and a polynomial mapping was later added to the B-Splines MPM [33, 23] resulting in an angular momentum-conserving model. Other studies explored least squares techniques [62] enabling the recovery of the affine/polynomial splines MPM [31], the effect of spatial and temporal discretisation errors, and the application of symplectic integrators [5]. In light of these developments, MPM represents an effective method for the simulation of a wide range of materials and phenomena.
However, in spite of recent implementations for computer graphics applications [24, 61], MPM implementations on GPUs have not yet been applied to compressible fluid dynamics. In fact, even when addressing gas dynamics and solids mechanics, MPM has been used only for the solids, while the compressible fluid dynamic problem has been solved through a second order WENO scheme [12]; moreover, such implementation is run on CPUs. Existing studies on MPM for compressible flows focused on numerical convergence and error analysis [57] and on modelling and expanding the range of applications for the method [65, 50]. Nonetheless, to the best of the authors’ knowledge, no studies exist that are dedicated to the GPU based, High Performance Computing aspects of MPM implementations for strongly compressible fluid dynamics problems.
Therefore, existing MPM implementations either present challenges in adapting to modern hardware or are not widely studied and employed in compressible fluid dynamics. A good example is given by high-order WENO and ENO schemes, that guarantee numerically excellent results in gas dynamics. For these schemes to work properly, the stencil should be able to grow in any direction into neighbouring mesh elements. From a computational point of view, this works very well on structured grids. However, designing proper structured grids for FSI is significantly time consuming. On the other hand, if unstructured grids are chosen, growing stencils lead to a large increase of halo transfers in communication among parallel processors, thus requiring a large consumption of memory bandwidth and generating a large amount of indirection, degrading the suitability of the methods for modern parallel hardware. Moreover, such schemes have to be coupled with appropriate methods for structural mechanics, that have to be suitable for modern parallel hardware too.
In this realm, MPM presents very interesting features as complete FSI solver, indeed:
  • it benefits from the choice of employing structured Cartesian grids, since continua discretisation is delegated to Lagrangian material particles, allowing the modelling of large deformations as well;
  • it does not require coupling among solvers, as it is suitable for structural mechanics problems as well.
However, even if MPM stemmed from fluid solvers, there are only a few studies addressing compressible fluid dynamics with MPM. In addition, the existing studies, some of which (e.g. [65]) predate work on numerical linear algebra on GPUs (e.g. [6, 35]), do not deal with parallelisation on modern hardware. In addition, the MPM community has preferentially paid attention to continuum mechanics modelling.
With this work we aim to address only the fluid portion of a future FSI solver, i.e., the portion which is usually left to Eulerian CFD solvers. We focus on algorithm and data structures suitable for implementation on various parallel architectures, with a particular attention to HPC GPUs, and we address the problem of reducing limitations in treating obstacles that derives from the choice of a structured Cartesian grid, by means of a signed distance function approach.
The current study focuses on linear MPM to better study its scalability features, leaving the update to a higher order MPM for future work. However, we anticipate a discussion on how higher regularity base functions could fit our approach and how the parallel efficiency would be influenced by the extension of base function support on more elements. From the modelling point of view, to give a first assessment of MPM in simulating efficiently compressible gases on GPUs, we treat the gas as inviscid fluid, neglecting the effects of turbulence, and adopt an artificial viscosity based shock capturing technique.
In previous work [2] we ported an existing MPM algorithm [65, 64, 14] to CUDA-C [41], adapting it to GPUs architecture and obtaining promising speed-ups with respect to a C++ CPU version. The CUDA-C programming model was chosen because of its known performance capabilities on NVIDIA A100 GPUs [42, 9]. However, the model had shortcomings in its low-level character and specificity to the underlying architecture. In addition, while it can be made to work on, e.g., AMD GPUs via HIP with minimal changes, it cannot be ported to different architectures such as CPUs without substantial code modifications.
One of NVIDIA’s solutions for performance portability involves its own C++ compiler nvc++ [21, 43], using which one rewrites algorithm steps relying on C++ Standard Template Library (STL), specifying a parallel execution policy. To give an example, in this framework it is possible to parallelise the following for loop:
1 for (index_t i = 0; i < n; ++i)
2 {do some operations}
into a call to a for_each with a proper lambda or functor, e.g.:
1 std::for_each (std::execution::par, first_counting_iterator, last_counting_iterator, [](index_t i){do some operations});
Although the approach does not prevent the developer from introducing data-races or deadlocks, the code can be written only once, and then, depending on the option specified during compilation, run in parallel on multi-core CPUs or GPUs.
However, nvc++ high-level features come at a price, as the developer loses the control on memory handling and, in particular, on when Host-to-Device and Device-to-Host memory copies are performed. The code runs on host, any invocation to an algorithm with a parallel execution policy expresses the developer’s preference to run that algorithm in parallel. If compiled for GPUs, memory is copied to the device, computations are performed and then data is copied back to host. In physics and in engineering dynamics simulations, where each time step depends on the previous one, this means that the maximum reachable level of parallelisation corresponds to a single call to a parallel algorithm inside a traditional for loop for time advancing. Since memory fetches can be significantly more costly than computations, this is sub-optimal.
The present contribution expands the scope of [2] and overcomes the performance issues with nvc++ by designing a portable MPM code for transonic and supersonic compressible gas dynamics based on Thrust [44], a C++ template library for CUDA based on the Standard Template Library (STL). In addition to nvc++ features, Thrust introduces two types of STL like vectors, one for the host (typically, the CPU) and one for the device (typically, the GPU); this allows the user to specify when memory should be copied via explicit calls. For example, it’s possible to perform a single Host to Device copy when reading the initial data file, and one Device To Host copy whenever saving partial results on output data files, strongly reducing execution time related to memory fetches.
Moreover, as additional portability-enhancing features included in the implementation:
  • Thrust syntax is mostly compatible with STL syntax, adding just a few define preprocessor macros or templates. Thanks to this, almost the same source code can be compiled for parallel execution on GPUs with NVIDIA CUDA compiler nvcc, on AMD GPUs with the AMD compiler for ROCm/rocThrust, and on CPUs with g++ and clang++ compilers (when it is linked with a parallel library such as TBB, Intel’s Threading Building Blocks).
[…]
Notwithstanding the portability objectives, we designed the code for optimisation on modern graphics hardware accelerators and made further modifications to the original algorithm [14, 65]. For example, we:
  • added treatment of obstacles in gas dynamics simulations,
  • added a particle re-ordering step,
  • postponed particles movement to the end of the time loop,
and thus values of basis function evaluated at previous particle positions don’t need to be stored and can be computed on the fly instead, which is cheaper on modern hardware accelerators.
Specific development choices are described that enable optimisation on modern GPU-based HPC architectures and adaptability to different hardware, as well as generality of the approach for the purposes of its portability also towards different software. The accuracy and performance of the implementation are evaluated on a number of compressible flow benchmarks and results are compared with available approaches in the literature.
The paper is structured as follows. Section 2 contains the analytical and numerical formulation of the method. Section 3 describes the chosen algorithm and the current and implementation choices. Section 4 illustrates the results obtained on benchmark test cases, and the final section 5 contains a discussion of the results, draws the conclusions of the performed work, and prospects future developments.
2 Model Formulation
2.1 Model Equations and Weak Formulation
As mentioned, MPM takes advantages of both the Eulerian and the Lagrangian perspective; specifically, the Lagrangian particles discretisation approach affords flexibility in modelling various kind of phenomena and materials, while the Eulerian grid discretisation is beneficial for scalability as the grid allows the mediation of particle-particle interactions. For the mathematical modelling we choose the Lagrangian framework, and we model compressible fluids in a domain $\Omega$ governed by the conservation laws for mass, linear momentum and energy densities [27, 64, 19]:
$\begin{cases}\dfrac{d\rho}{dt}+\rho\nabla\cdot\mathbf{v}=0\\[5.69054pt] \rho\dfrac{d\mathbf{v}}{dt}=\nabla\cdot\bm{\sigma}+\rho\mathbf{b}\\[5.69054pt] \rho\dfrac{de}{dt}=\bm{\sigma}:\bm{\dot{\varepsilon}}-\nabla\cdot\mathbf{q}+\rho s\end{cases}$ (1)
where $\rho$, $\mathbf{v}$, $\bm{\sigma}$, $\rho\mathbf{b}$, $p$, $e$, $\bm{\dot{\varepsilon}}$, $\mathbf{q}$ and $s$ represent respectively the mass density, the velocity, the stress tensor, the external forces density, the pressure, the specific internal energy, the strain rate tensor of the fluid, the heat flux and the heat source.
In this work, we:
  • enforce mass conservation by keeping constant the number of Lagrangian material points (particles) along the simulations, as well as each particle mass,
  • choose to model the fluid as a bi-atomic perfect gas, governed by the state equation
    $p=(\gamma-1)\rho e$ (2)
    where $e$ is the specific internal energy and $\gamma$ is the ratio between specific heat at constant pressure and constant volume, $\gamma=c_{p}/c_{v}=1.4$ for a bi-atomic gas.
To obtain the final discretisation of the equations, a weak form of the force balance expressed by the motion equation is derived [51, 50]. Having chosen a proper functional space $V$ of regular enough functions, we multiply the momentum conservation equation in (1) by an arbitrary $\mathbf{w}\in V$, integrate over the fluid domain $\Omega$ and integrate by parts the divergence term.
The weak formulation of the problem then reads:
Find $\mathbf{v}$ such that, $\forall\mathbf{w}\in V$
[…]
where $\mathbf{\dot{v}}$ denotes the total, or Lagrangian, velocity derivative, $\mathbf{w}$ is a generic test function and $\bm{\tau}$ is the stress on the surface. In the following, we simplify the model assuming the fluid to be inviscid, and we neglect external forces, heat sources and heat flux. We thus get the Euler equations for an inviscid, compressible fluid:
$\begin{cases}\dfrac{d\rho}{dt}+\rho\nabla\cdot\mathbf{v}=0\\[5.69054pt] \rho\dfrac{d\mathbf{v}}{dt}=-\nabla p\\[5.69054pt] \rho\dfrac{de}{dt}=-p\nabla\cdot\mathbf{v}\end{cases}$ (4)
Note that, in the above expression, the fluid stress tensor is simplified from:
$\sigma_{ij}=2\mu\dot{\varepsilon}_{ij}-\tfrac{2}{3}\mu\varepsilon_{kk}\delta_{ij}-p\delta_{ij}$
to:
$\sigma_{ij}=-p\delta_{ij}$
which greatly simplifies the energy conservation equation as well as the momentum equation.
Finally, following [63], we make use of a combination of von Neumann-Richtmyer [59] and Landshoff [36] artificial viscosity as shock-capturing scheme, relying on numerical damping to handle numerical oscillations in the vicinity of discontinuities, and computing them as part of the solution to problem (4). Specifically, application of this artificial viscosity-based shock-capturing method involves the addition of an artificial term $q$ to pressure $p$, substituting the pressure $p$ to the sum $p+q$ in problem (4). The artificial term is added only in compression regions, identified by the local velocity divergence being less than zero. Its expression reads
$q=\begin{cases}c_{0}\left(\rho\,h\,\mathrm{div}(v)\right)^{2}-c_{1}\,h\,\sqrt{\gamma\tfrac{p}{\rho}}\,\mathrm{div}(v)\,\,&\mathrm{div}(v)<0\\ 0\quad&\mathrm{div}(v)\geq 0\end{cases}$ (5)
where $c_{0},c_{1}$ are dimensionless constants and $h$ is the characteristic length of a grid cell. The effects of the two components of $q$ are shown in section 4.
This dynamically comprehensive yet relatively simple equation set is chosen as it enables testing the performance portability of the MPM implementation on compressible gas dynamics benchmarks [56, 47] and comparison with existing approaches [65, 50].
The continuity and energy conservation equations will be directly discretised and used to update particle properties in the algorithm (see section 3 and algorithm 1), while the discretisation of the weak form of the motion equation is related to the link between the Eulerian grid and the Lagrangian particles, and is treated in more depth in the next section.
2.2 Numerical Method
The link between the Eulerian grid and the Lagrangian particles is given by the Particle-to-Grid (P2G) and Grid-to-Particle (G2P) mappings, which are the building blocks of the computational algorithm, and that can be derived during the discretisation of the weak form (3), that leads to a semi-discrete equation of the form [^1] Here and in the following, when it is useful to distinguish, grid quantities are written in lower case Roman characters, while particle ones are written in either upper case Roman or Greek characters; moreover, $i$ denotes grid nodes index, while $k$ denotes particles index.: $\forall i=0\ldots N_{h}-1\quad m_{i}\mathbf{\dot{v}}_{i}=\mathbf{f}_{i}$ (6) where $N_{h}$ denotes the number of grid nodes.
The latter can be obtained following the standard MPM procedure (see e.g. [51]), including five main steps: 1. discretise the material domain into a mesh of material sub-domains, with mass concentrated at material points $\rho(\mathbf{x},t)=\sum_{k=0}^{N_{p}-1}M_{k}\delta(\mathbf{x}-\mathbf{X}_{k})$ (7) where $N_{p}$ denotes the total number of material points, and insert (7) into (3); 2. discretise space with a standard Finite Elements Method (FEM) grid; 3. approximate any field and test function relying on FE grid shape functions $u_{i}$; 4. apply arbitrariness of test function to get the semi-discrete equation; 5. possibly, lump the mass matrix $m_{i}\!=\!\sum_{j}m_{ij}\!=\!\sum_{j}\sum_{k}M_{k}u_{i}(\mathbf{X}_{k})u_{j}(\mathbf{X}_{k})\!=\!\sum_{k}M_{k}u_{i}(\mathbf{X}_{k})$ (8) where $u_{i}(\mathbf{X}_{k})$ denotes the basis function of the solution space and test functions space centred on the $i$-th grid node and evaluated in the $k$-th particle position, $\mathbf{X}_{k}=(X_{k},Y_{k})$.
It is worth noting that it is possible to derive the same formulation also by adopting the standard Galerkin-Finite Element procedure, provided material points are considered as moving quadrature points [19]. In fact, standard MPM can be seen as a modification of traditional FEM, in which the integration points are not related to the Finite Element mesh and are Lagrangian, and thus suitable for history dependent models too [14].
Following this method, we start from the motion equation expressed in weak formulation, which reads “find $\mathbf{v}\in V$ such that equation (3) holds $\forall\mathbf{w}\in V$”, $V$ being a proper Hilbert space. To discretise the problem, we introduce a finite dimensional sub-space $V_{h}=span\{u_{i}(x,y)\}$ , project equation (3) onto $V_{h}$, and, since it has to hold for any $\mathbf{w}_{h}\in V_{h}$, we find $\mathbf{v}_{h}$ such that equation (3) holds for each function in $V_{h}$ base.
Since $V_{h}$ is finite dimensional, it is possible to express any function in terms of finite sum, e.g.
$v_{h}(x,y)=\sum_{i}v_{i}u_{i}(x,y)$ (9)
In other words, we write equation (3) choosing as test function the base functions $u_{i}$ centred on the $i-th$ grid node, and we solve numerically the integrals by means of quadrature rules, evaluating the integral at a finite set of points, called the integration points $p$, and perform the weighted sum
[…]
If the material points are taken as integration points, that is, if basis functions are evaluated in the $k-th$ particle position, with weights being their volumes (i.e. $\omega_{p}=\tfrac{M_{k}}{\rho_{k}}$), then FEM equations reduces to MPM ones. In particular,
$\mathbf{f}^{internal}_{i}=-\int_{\Omega}\bm{\sigma}\nabla u_{i}d\Omega\simeq-\sum_{k=0}^{N_{p}-1}\bm{\sigma}_{k}\tfrac{M_{k}}{\rho_{k}}\nabla u_{i}(X_{k},Y_{k})$ (11)
which is the internal component of the force in the semi-discrete equation (6), where we have already applied mass matrix lumping.
In particular, equation (11) expresses the Particle-to-Grid mapping of internal forces, which, in our inviscid gas model with $\bm{\sigma}=-p\mathbb{I}$, maps the pressure force from the material points to the grid nodes. To distinguish this from the Particle-to-Grid which does not involve moving the derivative to the test function, we call operations as (11) P2GD.
P2GD is one of the four building blocks of MPM algorithm, the others being:
  • P2G, without weak derivatives, which is needed since grid variables are reset at every time step (see sec. 3);
  • G2P, which maps quantities back to the particles after equations have been solved numerically;
  • G2PD, which also computes in-place a weak derivative.
In this work, we define these operations, which we call kernels following GPU literature, in a general way to underline their computational structure for later implementation and performance analysis, even though they are applied only as written in the algorithm in section 3. More details on how these operations are derived can be found in the original papers for the standard MPM [51], in [64, 65] for a first application to compressible fluids, and in [14, 63, 19] for later reviews.
Denote by $\mathbf{a}=[a_{i}]\in\mathbb{R}^{N_{h}}$ the array of values of some quantity $a(x,y)$ at the grid vertices $(x_{i},y_{i}),i=0\ldots N_{h}-1$, and let $\mathbf{A}=[A_{k}]\in\mathbb{R}^{N_{p}}$ be the array of values of some quantity $A(X,Y)$ at the material points $(X_{k},Y_{k}),k=0\ldots N_{p}-1$.
Particle-to-Grid and Grid-to-Particle mappings read:
$a_{i}=\mathrm{P2G}(\mathbf{A};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}A_{k}\ u_{i}(X_{k},Y_{k}),\quad i=0\ldots N_{h}-1$ (12)
$A_{k}=\mathrm{G2P}(\mathbf{a};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}a_{i}\ u_{i}(X_{k},Y_{k}),\quad k=0\ldots N_{p}-1$ (13)
$\left(\nabla a\right)_{i}=\mathrm{P2GD}(\mathbf{A};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}A_{k}\ \left(\nabla u\right)_{i}(X_{k},Y_{k}),\quad i=0\ldots N_{h}-1$ (14)
$\left(\nabla A\right)_{k}=\mathrm{G2PD}(\mathbf{a};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}a_{i}\ \left(\nabla u\right)_{i}(X_{k},Y_{K}),\quad k=0\ldots N_{p}-1.$ (15)
It is useful to generalise these kernels to the case where the mapped quantity is not stored in memory but computed on the fly as a function of other state variables, as in equation (11). To this end, let $\mathcal{F}:\mathbb{R}^{m}\to\mathbb{R}$, and rewrite the above kernels as follows :
$a_{i}=\mathrm{P2G}(\mathcal{F},\mathbf{A}^{(0)},\ldots,\mathbf{A}^{(m-1)};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}\mathcal{F}\left({A}_{k}^{(0)},\ldots{A}_{k}^{(m-1)}\right)\ u_{i}(X_{k},Y_{k}),\quad i=0\ldots N_{h}-1$ (15)
$A_{k}=\mathrm{G2P}(\mathcal{F},\mathbf{a}^{(0)},\ldots,\mathbf{a}^{(m-1)};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}\mathcal{F}\left({a}_{i}^{(0)},\ldots{a}_{i}^{(m-1)}\right)\ u_{i}(X_{k},Y_{k}),\;k=0\ldots N_{p}-1$ (16)
$\left(\nabla a\right)_{i}=\mathrm{P2GD}(\mathcal{F},\mathbf{A}^{(0)},\ldots,\mathbf{A}^{(m-1)};\ x_{i},y_{i}):=\\ \sum_{k=0}^{Np-1}\mathcal{F}\left({A}_{k}^{(0)},\ldots,{A}_{k}^{(m-1)}\right)\ \left(\nabla u\right)_{i}(X_{k},Y_{k}),\\ i=0\ldots N_{h}-1$ (17)
$\left(\nabla A\right)_{k}=\mathrm{G2PD}(\mathcal{F},\mathbf{a}^{(0)},\ldots,\mathbf{a}^{(m-1)};\ X_{k},Y_{k}):=\\ \sum_{i=0}^{N_{h}-1}\mathcal{F}\left({a}_{i}^{(0)},\ldots{a}_{i}^{(m-1)}\right)\ \left(\nabla u\right)_{i}(X_{k},Y_{K}),\\ k=0\ldots N_{p}-1.$ (18)
For the numerical discretisation of the Eulerian aspects of the method, we choose a Cartesian grid of quadrilateral elements. Among the possible choices, this one lends itself naturally to computationally efficient refinements through a quad-tree approach [46]. Moreover, we opt for Lagrangian hat functions as basis for the functional spaces of both the solution and the test function, setting our treatment in a $\mathbb{Q}_{1}$ framework. [^2] In sec. 5 we discuss the extension to higher order basis functions.
This approach corresponds to bi-linear weighting function of standard MPM, and is known to cause cell-crossing error and spurious oscillations in the computed quantities. The reason can be easily understood by the above equations: when $u$ is (bi)-linear, $\nabla u$ is constant, but potentially different in neighbouring cells. Thus, when crossing cells, particles might experience discontinuous oscillations. This issue can be numerically tackled with many strategies, among which adding artificial viscosity [63] or choosing higher order basis functions. For simplicity of implementation, we choose the first, leaving the second for a future work.
With respect to standard MPM [64, 14], the present algorithm adds a step to take into account the flow around obstacles. Namely, we subtract the inward component of both linear momentum and force from the grid nodes whose cell is marked as an obstacle cell in following equations (stemming from the method for granular materials [3]):
$(m\mathbf{v}_{i}^{o})^{\prime}=m\mathbf{v}_{i}^{o}-\dfrac{m\mathbf{v}_{i}^{o}\cdot\mathbf{n}_{i}^{o}-|m\mathbf{v}_{i}^{o}\cdot\mathbf{n}_{i}^{o}|}{2}\mathbf{n}_{i}^{o}$ (19)
$(\mathbf{f}_{i}^{o})^{\prime}=\mathbf{f}_{i}^{o}-\dfrac{\mathbf{f}_{i}^{o}\cdot\mathbf{n}_{i}^{o}-|\mathbf{f}_{i}^{o}\cdot\mathbf{n}_{i}^{o}|}{2}\mathbf{n}_{i}^{o}$ (20)
where the subscript denotes the $i-th$ grid node, the superscript $o$ stands for obstacle and $\mathbf{n}$ is the normal unit vector, positive if directed outward from the obstacle.
Normal unit vectors to a domain profile $\partial\Omega$ are computed as gradients of the Signed Distance Function, defined as
$sd(\mathbf{x})=\begin{cases}-\inf_{y\in\partial\Omega}d(\mathbf{x},\mathbf{y})\quad\mathbf{x}\in\Omega\\ \inf_{y\in\partial\Omega}d(\mathbf{x},\mathbf{y})\quad\mathbf{x}\notin\Omega\end{cases}$ (21)
where $d(x,y)$ is the Euclidean distance. Taking the gradient yields [^3] An algorithm for signed distance function computation can be found in [10], and an example in a machine learning based framework that implicitly represents geometries from noisy surface points can be found in [13].
$\nabla sd(\mathbf{x})=\mathbf{n}(\mathbf{x})\quad\mathbf{x}\in\partial\Omega$ (22)
Finally, we use explicit time integration [14], whose stability is subject to a CFL criterion
$\Delta t<\alpha\dfrac{\Delta x_{min}}{v_{max}}$ (23)
where $\alpha$ collects constants. The range of admissible time steps is bounded from above by the ratio between the smallest grid cell dimension $\Delta x_{min}$ and the highest speed $v_{max}$, considering both particles velocity and speed of sound $c_{s}=\sqrt{\gamma\tfrac{p}{\rho}}$. The efficiency cost of enforcing small time steps is balanced by the need of resolving fast dynamics. In addition, explicit-in-time methods sidestep the need to solve global linear systems at each time step, thus enjoying better scalability properties than implicit methods.
3 Portable MPM Implementation
This section describes the algorithm that was presented in section 2 and how it was implemented to ensure portability on different architectures and efficiency on GPUs. The development of the code benefited in part from publicly available Open Source libraries: Thrust [^4] https://github.com/NVIDIA/cccl/tree/main/thrust, JSON [^5] https://github.com/nlohmann/json, and quadgrid [^6] https://github.com/carlodefalco/quadgrid. In detail, the algorithm implementation is composed of the following steps:
  1. 1.
    Data generation: the creation of simulations input is performed separately from code execution. For simple geometries it is handled via ad-hoc C++ scripts; for complex geometries, requiring a careful computation of the signed distance function, an Octave script is used instead, to take advantage of Octave built-in functions. In both cases, since objects are considered at rest, normal unit vectors are computed only once, and boundary cells are marked with a flag corresponding to the boundary condition. Finally, the initial data file is saved in JSON format.
  2. 2.
    MPM class object construction: since the most computationally costly operations are the memory related ones, such as I/O and copies between host and device, a class object for the method is constructed only once, reading grid and particles data from the input file, data is saved on host memory and copied to device memory. So doing, is possible to construct functor objects which will be called on the device during the time loop, without the need of further copies.
  3. 3.
    The time advancing loop starts; inner steps, except from output operation, are implemented as Thrust or STL for _eachs, depending on the architecture chosen at compile time.
    1. (i)
      Grid variables are reset and current grid variables are overwritten with the ones saved during construction stage, performing a device to device copy. This step can be performed since all history dependence is maintained by Lagrangian particles.
    2. (ii)
      Equations (12) and (17) are applied to map particles mass, momentum and force to the grid; during force P2GD, artificial viscosity is applied as in (5).
    3. (iii)
      Boundary conditions are enforced on grid nodes momentum and force. Specifically, at each time step and for each boundary node marked as of obstacle type, inward components of momentum and force are subtracted as in equations (19) and (20).
    4. (iv)
      Grid nodes momentum is advanced in time.
    5. (v)
      Velocity and acceleration generalised G2P (16) is applied, and next, velocity is updated. Acceleration itself is not stored, but computed as the result coming from a local G2P of the relevant nodes ratio between force and mass; moreover, G2P kernels are performed on particles index, as shown in the snippet in Fig. 1, since this has proved to provide some acceleration with respect to an approach based on grid nodes index. Following standard MPM [14], we store the G2P velocity in a temporary velocity vector $\mathbf{\overline{V}}$ and don’t use it to update particles velocity; however, we postpone its usage.
    6. (vi)
      Momentum P2G (12) is performed.
    7. (vii)
      Boundary conditions on the updated grid values are enforced.
    8. (viii)
      Particles strain rate tensor (velocity gradient) is computed as the G2PD (18) of the ratio between grid nodes momentum and mass, and then particles properties are updated; also in this case, the particle index is used for the G2PD kernel, and artificial viscosity is applied in the update of internal energy density.
    9. (ix)
      Particles position are updated using $\mathbf{\overline{V}}$. Thus, the time advancing scheme is the same of [14]. However, particles moving is postponed, so that, in all the above steps, basis functions and their gradient can be computed on the fly, without the need to store the value they had at former particles position.
    10. (x)
      Particles-to-grid connectivity map is updated with new positions - this step is computationally efficient also thanks to the choice of a Cartesian grid.
    11. (xi)
      Only if the initial data options ask for $N$ intermediate outputs, a device to host memory copy and an output from host to disk are performed at $N$ moments equidistant in time.
    12. (xii)
      Optionally, a re-ordering algorithm is applied to keep data locality, i.e. to keep contiguous particles in the physical simulation also close in memory. The algorithm is based on an indexing vector built iterating over the structure that stores information about which particles are contained in every grid cell, such that the n-th element of the ordering vector indicates the current sorting particle index which has to be at n-th position after re-ordering.
    13. (xiii)
      A control on the CFL condition (23) is enforced, taking into account particles velocity (momentum over mass and $\mathbf{\overline{V}}$) and the speed of sound.
    14. (xiv)
      Time step is updated, loop ends.
  4. 4.
    A final device-to-host memory copy and output is made.
Algorithm 1 summarises the for loop procedure.
Algorithm 1 Time loop sketch
1:while  $t<t_{f}$ do
2:     Grid reset
3:     Mass, momentum and force P2G
$\qquad m=\mathrm{P2G}(M),\ m\mathbf{v}=\mathrm{P2G}(M\mathbf{V}),\ \mathbf{f}=\mathrm{P2GD}((P+Q)\tfrac{M}{\rho})$
4:     Boundary conditions enforcement on $m\mathbf{v},\,\mathbf{f}$ obstacle treatment as in (19), (20)
5:     Momentum equation solution on the grid
$\qquad(m\mathbf{v})^{\prime}=m\mathbf{v}+\Delta t\mathbf{f}$
6:     G2P mapping
$\qquad\overline{\mathbf{V}}=\mathrm{G2P}(\tfrac{m\mathbf{v}}{m}),\ {\mathbf{A}}=\mathrm{G2P}(\tfrac{\mathbf{f}}{m})$
7:     Velocity update
$\qquad\mathbf{V}^{\prime}=\mathbf{V}+\Delta t\mathbf{A}$
8:     Momentum P2G
$\qquad m\mathbf{v}=\mathrm{P2G}(M\mathbf{V})$
9:     Boundary condition enforcement on $m\mathbf{v}$
10:     Particles properties update:
$\qquad\nabla\mathbf{V}=\mathrm{G2PD}(\tfrac{m\mathbf{v}}{m})$
$\qquad E^{\prime}=E+\Delta t\dfrac{-(P+Q)}{\rho}\nabla\cdot\mathbf{V}$
$\qquad\rho^{\prime}=\dfrac{\rho}{1+\Delta t\nabla\cdot\mathbf{V}}$
$\qquad P^{\prime}=(\gamma-1)\rho^{\prime}E^{\prime}$
11:     Particles moving $\mathbf{X}^{\prime}=\mathbf{X}+\Delta t\overline{\mathbf{V}}$
12:     Optional particles re-ordering
13:     CFL condition (23) check & time advance
14:end while
This algorithm is of special interest for acceleration on GPUs since it can be written in terms of the four kernels (15)-(18), which present interesting parallelisation potentialities. Grid-to-Particles operations are intrinsically Single Instruction Multiple Data (SIMD) parallelisable, since data is projected from grid nodes to particles, and thus particles are independent from one another during writing stage. Particles-to-Grid kernels present the same potentialities, however, if they are implemented strictly following their analytical definition, they present a data race: different particles need access in both read and write mode to the same grid nodes.
In this work, we chose to patch the Particles-to-Grid and Grid-to-Particles kernels written in the external quadgrid library to adapt them to GPU execution via Thrust rather than writing ad hoc versions in the code, at the expense of some further optimisation that could have been added by exploiting the peculiarities of our test cases, to arrive at more general and portable conclusions. In this way, there is only one computational kernel per operation (15)-(18). In any case, preliminary tests have shown that applying a kernel-fusion approach [44] where possible wouldn’t have affected the execution time order of magnitude.
To better illustrate the concepts introduced in the last paragraph, we sketch and discuss below the implementation and use of the two main kernels in the algorithm, namely the Grid-to-Particle (G2P) operation (13) and the Particle-to-Grid operation (12). A simplified sketch of the implementation of the former is given in Fig. 1, while the latter is described in Fig. 4; the reported code corresponds roughly to the actual implementation available in the quadgrid library except for a few details that are left out only to avoid distraction as they are not interesting for the current discussion. As already noted, the G2P operation is easily implemented via a loop over all particles and lends itself quite naturally to SIMD parallelisation as all write operations are independent and may be carried out concurrently. The actual looping is performed via the for_each method, which can be used to perform any operation on any number of variables associated to a particle by using a counting iterator and a suitably constructed callable object as shown in Fig. 3 and Fig. 2. The P2G operation is implemented along lines very similar to those followed for G2P. Specifically, as shown in Fig. 4 the looping is done over the list of particles just as for G2P, which is made possible by the index array ptcl_to_grid that lists, for each particle, the grid cell within which it is located. Efficiently building the latter data structure heavily depends on the use of structured grid. The main difference between P2G and G2P consists in the need, in the former, to work around a data race that is created when threads working on different particles try to write data related to the same neighbouring cell vertex. The simplest approach, followed in this work, is to make the write operations atomic by using atomicAdd (as shown for sake of example in Fig. 4), which is readily available in CUDA but can also be re-implemented in standard C++ based on the std::atomic_ref class template introduced in the C++20 standard. Although our tests show that the use of atomics does not seem to lead to much of a performance degradation on GPUs, we discuss in a later section a possible alternative based on data reordering and mesh colouring that can be used to avoid atomics.
1 template<class callable_t>
2 G2P (idx_t iptcl, callable_t *grid_variable,
3 ptcl_variable_t *ptcl_variable) {
4
5 idx_t icell = ptcl_to_grid[iptcl];
6 real_t x = ptcl_x[iptcl];
7 real_t y = ptcl_y[iptcl];
8
9 for (idx_t inode = 0; inode < 4; ++inode) {
10 real_t N = shape_function (x, y, inode, icell);
11 idx_t ginode = local_to_global (icell, inode);
12
13 ptcl_variable[iptcl] +=
14 N * (*grid_variable)(ginode);
15 }
16
17 };
Figure 1: Example template function to demonstrate the approach used to implement the G2P kernel. This is not the actual code used but a conceptualised, reduced example to explain the basic idea. In particular note that in this simplified snippet (and some of the following) we use, for sake of simplicity of presentation, standard pointers, while the actual implementation uses some form of iterator.
1 struct x_velocity_t {
2 grid_data_t *gdata = &grid_data;
3 real_t operator()(idx_t ginode) {
4 real_t vx = gdata -> get (ginode, "x_momentum");
5 real_t m = gdata -> get (ginode, "mass");
6 return vx / m;
7 }
8 } *x_velocity;
9
10 x_velocity_g2p =
11 [x_velocity, ptcl_x_velocity] (idx_t iptcl) {
12 G2P (iptcl, x_velocity, ptcl_x_velocity);
13 };
Figure 2: Instantiation of a callable object that implements grid to particle interpolation of the $x$–component of the velocity. This is not the actual code used but a conceptualised, reduced example to explain the basic idea. In the code above, the purpose of the callable object pointed to by x_velocity is to compute on–the–fly a function of the state variables at a grid vertex, so that its value can be interpolated at the particle locations without requiring extra storage or data transfers.
1 counting_iterator<idx_t> first_p(0);
2 ounting_iterator<idx_t> last_p(num_particles);
3 for_each (execution_policy,
4 first_p, last_p, x_velocity);
Figure 3: Example of performing an actual G2P operation by iterating over all particles by means of the for_each method and counting iterators.
1 template<class callable_t>
2 P2G (idx_t iptcl, ptcl_variable_t *grid_variable,
3 callable_t *ptcl_variable,) {
4
5 idx_t icell = ptcl_to_grid[iptcl];
6 real_t x = ptcl_x[iptcl];
7 real_t y = ptcl_y[iptcl];
8
9 for (idx_t inode = 0; inode < 4; ++inode) {
10 real_t N = shape_function (x, y, inode, icell);
11 idx_t ginode = local_to_global (icell, inode);
12
13 atomicAdd (grid_variable + iptcl,
14 N * (*ptcl_variable)(ginode));
15 }
16
17 };
Figure 4: Example template function to demonstrate the approach used to implement the G2P kernel. This is not the actual code used but a conceptualised, reduced example to explain the basic idea. In particular note that in this simplified snippet (and some of the following) we use, for sake of simplicity of presentation, standard pointers, while the actual implementation uses some form of iterator.
4 Benchmarks
In this section we show the results of 2D numerical simulations of gas dynamics obtained with the described implementation of MPM, with a focus on transonic and supersonic gas flow over solid obstacles. A first general evaluation of accuracy in subsection 4.1 is followed by an assessment of the sensitivity of the results to the choice of numerical parameters in subsection 4.2; in particular, the effects of numerical viscosity components on mass density are shown for a test case with analytical solution in subsection 4.2.
In subsection 4.3 we then test the efficiency and performance of the implementation in providing parallel scalability on NVIDIA A100 GPUs. Lastly, in subsection 4.4, we assess the portability of the developed code to different CPUs and GPUs architecture, which is our main claim.
The experiments were carried out on NVIDIA A100 SXM4 40 GB GPUs, Intel Xeon Platinum 8260 and AMD EPYC Rome 7402 CPUs at Leonardo S.p.A.’s in-house HPC facility davinci-1, and on AMD EPYC 7313, AMD EPYC 7413 CPUs and AMD Radeon Instinct MI210 GPUs at MOX Laboratory for Modeling and Scientific Computing at Politecnico di Milano’s Mathematics Department.
All tests are performed in double precision, i.e. using FP64. Depending on the architecture, better performance could be achieved with lower precision floating points; for example, the NVIDIA GPUs here employed, which date back to 2020, should gain a 2x factor in computational time when working with FP32. However, despite the code supports different choices, we have focused only on doubles, since floats could lead to numerical difficulties in some test cases, and already the slightly more recent AMD MI 210 GPUs (release date December 2021) are able to handle double and single precision floats without any performance degradation.
4.1 Test cases
In the following numerical tests, we adopt non-dimensional units as in [47], and we model the fluid as a perfect, bi-atomic gas, with specific heat ratio:
$\gamma=\dfrac{c_{p}}{c_{v}}=1.4$ (24)
Moreover, except in the Sod Shock Tube, in the far field away from any obstacle, we set fluid density and pressure as:
  • $\rho_{\infty}=1.4$;
  • $p_{\infty}=1$;
where the subscript $\infty$ denotes unperturbed conditions. Thus, we obtain a unitary speed of sound:
$c_{s,\infty}=\sqrt{\gamma\ \dfrac{p_{\infty}}{\rho_{\infty}}}=1$ (25)
and an unperturbed Mach number equal to the reference speed:
$M_{\infty}=\tfrac{v_{\infty}}{c_{s,\infty}}=v_{\infty}$ (26)
For time integration, in our code we enforce the time step to be:
$\Delta t\leq\dfrac{1}{2}\dfrac{h_{min}}{v_{max}+c_{s,max}}$ (27)
This requirement is more conservative than necessary, and, in the limits of the CFL condition, could be partially relaxed. However, we choose it to guarantee that particles can’t move across more than one grid cell per time iteration; so doing, in the framework of our algorithm, we avoid the risk that particles pass through any boundary, and ensure that their motion will respect boundary conditions.
In all test cases except in the Sod Shock Tube, initial particles position is sampled by a uniform random distribution in the fluid domain. Indeed, a sufficient randomness of particles position, especially in particles recycling from outlet to inlet, is fundamental to obtain stable and physically meaningful simulations, in agreement with results in the literature on Particle-In-Cell methods in plasma physics. On longer runs (i.e. above tens of physical seconds) a symplectic time integration scheme might be useful [34].
In the following, we denote with $ppc$ the average number of fluid particles per fluid cell, with $N_{C}$ the total number of cells, and with $N_{P}$ the total number of particles.
Sod Shock Tube
In the Sod Shock Tube, two compressible fluids, initially at rest, are divided by a diaphragm. The first fluid has $\rho_{1}=1$ and $p_{1}=1$, the second $\rho_{2}=0.125$ and $p_{2}=0.1$. At time $t=0$ the diaphragm is removed, and the system evolution is studied; the following phenomena are expected:
  • the two initial zones shrink progressively towards domain borders;
  • a rarefaction wave propagates toward the high density zone;
  • in the position where the diaphragm was, which now is moving towards low densities, a contact discontinuity develops;
  • a shock wave propagates towards low density zones.
In order to reproduce the 1D problem, we introduce a mesh of $1\times 100$ cells; moreover, to be able to compare the MPM solution with the reference one, we also set initial particle position as in [65, 64].
In Figure 5 we show the final particles position, together with a comparison of the computed solution, its linear interpolation on the grid, and the exact solution computed following the procedure shown in [56]. Results are in very good agreement with the MPM literature [65, 19].
Figure 5: Sod Shock Tube: comparison of MPM solution computed on particles, interpolated on the grid, and exact solution computed as in [56]
Supersonic flow past solid obstacles
We then consider two cases of supersonic, Mach 3 flow:
  • (a)
    past a cylinder of radius $0.25$ and centre $(0.6,1)$ in a $[0,4]\times[0,2]$ rectangular domain; and
  • (b)
    past a rectangular-shaped step $[0.6,3]\times[0,0.2]$ in a $[0,3]\times[0,1]$ rectangular domain.
In both cases, we enforce wind tunnel–like boundary conditions, with inflow conditions at the left vertical boundary, outflow conditions at the right horizontal boundary, and slip conditions at the top and bottom boundaries.
Specifically, we set:
$\begin{cases}\mathbf{v}_{inflow}=(3,0)\\ \mathbf{v}_{top}\cdot\mathbf{n}=\mathbf{v}_{bottom}\cdot\mathbf{n}=0\end{cases}$
where $\mathbf{n}$ denotes the unit normal vector. A particles-recycling approach is adopted, whereby particles leaving the domain at the outflow boundary are regenerated in a random cell at the inflow boundary with inflow conditions, keeping the total number of particles constant throughout the simulation. In test case (a), as gas particles travel past the cylinder, a shock wave develops at the front and two oblique shocks in the rear. The first shock is reflected by walls and interacts with the back shocks; see Figure 6 and its caption for numerical details.
Figure 6: Mach 3 flow past a cylinder: pressure at $t=10\,s$, $N_{C}=10^{6}$, $ppc=16$, $N_{P}=15.6\cdot 10^{6}$
Figure 7: Mach 3 flow past a step: pressure at $t=12s$, $N_{C}=326.7\cdot 10^{3}$, $ppc=16$, $N_{P}=4.4\cdot 10^{6}$
In the step test (b), a frontal shock develops that is later reflected from the top wall and successively from the step see Figure 7 and its caption for numerical details.
The MPM implementation is able to correctly simulate the main flow features, in case of both round shapes and sharp edges and corners. Computed shock-waves angle and flow deflection angle (Figure 6) are found to be in agreement with experimental references [20] as well as with results with other codes, e.g., deal.II, when the same polynomial order is used for basis functions. In particular, the cylinder test case is simulated in the deal.II documentation [39] [^7] https://www.dealii.org/current/doxygen/deal.II/step_69.html#Results. With our implementation, we are able to reproduce the main features of the front shock starting from a very low resolution, with details, especially in the rear, getting better as resolution increases; more details are shown in subsection 4.2. Overall, the results are in agreement with [39], while the lower level of detail in comparison with some results in the literature ([16, 38, 27]) is likely due to the first order space discretisation used here.
Transonic flow past an aerofoil
Next, we test the method in simulating compressible, transonic flow past an OAT15a aerofoil at Mach 0.73, to assess the ability of the implementation in dealing with more complex-shaped obstacles. The signed-distance function approach, in conjunction with the implemented obstacles treatment (19),(20), naturally fits the aerofoil profile (Figure 8). Since, rather than using an analytic expression, the profile and its normal unit vectors are generated reading from a file containing a set of points, the method proves to be sufficiently general and not bounded to particular geometric shapes.
Within the $[0,0.4]\times[0,0.27]$ domain, we enforce inflow conditions at the left boundary, with
$\mathbf{v}_{\infty}=(0.73,0)$ (28)
and outflow conditions at other boundaries. To test the physical reliability of the approach, we choose a large attach number to emphasise dynamical effects and display velocity magnitude and streamlines (Figure 9).
These are found to be in line with what is expected from physical considerations: we can see airflow speeding up as it approaches the aerofoil, with a higher velocity magnitude in the top area with respect to the bottom one.
Figure 8: Signed distance function for OAT15a aerofoil at $\alpha=-3.5^{\circ}$
Figure 9: OAT15a Mach 0.73 flow: velocity streamlines at $\alpha=25^{\circ}$, $t=1s$, $N_{C}=247\cdot 10^{3}$, $ppc=9$, $N_{P}=2.18\cdot 10^{6}$.
Taylor-Green vortex
In the third test, we address the Taylor-Green Vortex test case, a well studied, fluid-only benchmark to assess both viscous and inviscid fluid dynamics codes in the literature, usually at subsonic regime ([11, 7, 54]); here we simulate the case in supersonic Mach 3 conditions.
In the inviscid case, the dynamics is strongly affected by problem symmetries, leading to “impermeable box boundary conditions” [7]. These symmetries are:
  • within the $[0,\pi]^{2}$ square-shaped computational domain, the fluid never crosses the lines $x=n\pi$ and $y=n\pi$; and
  • close to the boundaries, both the normal velocity $\mathbf{v_{\perp}}$ and the normal derivative ${\tfrac{\partial\mathbf{v_{//}}}{\partial\mathbf{n}}}$ of the tangential velocity $\mathbf{v_{//}}$ vanish.
We then set the initial conditions:
$\begin{cases}v_{x}(x,y)=v_{\infty}\sin\left(\tfrac{x}{\pi}\right)\cos\left(\tfrac{y}{\pi}\right)\\ v_{y}(x,y)=-v_{\infty}\cos\left(\tfrac{x}{\pi}\right)\sin\left(\tfrac{y}{\pi}\right)\\ p(x,y)=p_{\infty}+\tfrac{\rho_{\infty}}{16}\left(\left(\cos\left(\tfrac{2x}{\pi}\right)+\cos\left(\tfrac{2y}{\pi}\right)\right)\cdot 3-2\right).\end{cases}$
The numerical method captures the flow features with good accuracy as compared to results in the literature (Figure 10, cf. Figure 1b in [7]), confirming the general applicability of the MPM algorithm in simulating compressible fluid flow.
Figure 10: Taylor-Green vortex: computed velocity at $t=2s$, $N_{C}=3.94\cdot 10^{5}$, $ppc=9$, $N_{P}=3.55\cdot 10^{6}$
4.2 Sensitivity to numerical parameters
Artificial viscosity
Artificial viscosity introduced in section 2 is composed of two main terms: a quadratic one and a linear one. The first, introduced by von Neumann-Richtmyer [59], is a viscous term that, being added to the pressure in the vicinity of shock waves, reduces numerical oscillations and enables the capturing of the shock. The second, introduced by Landshoff [36], is a linear term added to the quadratic one to act specifically in the region behind the shock front. Further details on the application of this viscosity to MPM can be found in [63].
In this paragraph we study the effect of the artificial viscosity, comparing the final density obtained enforcing $c_{0}=1,c_{1}=1$ in equation (5) and shown in Figure 5, with the one obtained by:
  • turning off the artificial viscosity;
  • applying only the von Neumann-Richtmyer term, enhanced by a factor 10;
  • applying only the Landshoff term, enhanced by a factor 10.
Results are in agreement with theoretical expectations (Figure 11) and confirm the importance of this approach in the linear MPM scheme adopted here.
Figure 11: Sod Shock Tube - density trend at time $t=0.143$, varying artificial viscosity coefficients.
Resolution
Having addressed the applicability of the method to high-speed transonic and supersonic compressible gas dynamics, we focus on the cylinder test case and test the dependence of the results on key numerical parameters. Specifically, we vary the number of grid cells, $N_{C}$, and the average number of particles per cell, $ppc$, and evaluate the effect on the pressure of the flow.
First, sufficiently high spatial resolution must be used to actually show shock waves; indeed, notwithstanding the use of artificial viscosity as shock capturing technique, with $N_{C}\sim 10^{3}$ cells we are not able to see any shock. However, a run using $N_{C}\simeq 10^{4}$ cells is able to show flow features which are in agreement with what expected from the literature, as can be seen in Figure 12.
Besides resolution, supersonic compressible fluid dynamics requires further care in setting the average number of particles per cell. Indeed, $ppc$ affects stability in a twofold fashion.
First, for low $ppc$ values (e.g.,1), widespread oscillations are observed all along the simulation, due to almost randomly distributed empty cells. In a recent study, a $ppc$ value of 4 was found to be optimal for convergence of gas dynamics tests at lower speed [50], and, indeed, a $ppc=4$ solves this issue. However, for supersonic flow past a static obstacle, we observed that using $ppc=4$ might cause the formation of domain regions of empty or quasi-empty cells, which gets evident as grid resolution increases, during the initial transient phase from uniform, undisturbed conditions, as can be seen from Figure 13. Such emptying causes later formation of nonphysical cavitation-like effects during the transient that precedes equilibrium, which alter the dynamics, and, if particularly strong, can cause also spurious shock waves, that interact with the physical ones. These events can be prevented by increasing the average particles per cell number: at Mach 3, spurious shock waves become less relevant with $ppc=9$, and undetectable with $ppc=16$.
These effects, and in particular strong disruptions, are of no interest here, since the initial transient phase is non-physical, and derives only from the choice to start the simulations with undisturbed inflow conditions and particles position sampled from a uniform random distribution instead that from a previously computed equilibrium condition, e.g. the result of an already converged simulation. However, it could be a relevant aspect if fast changing dynamic situations have to be simulated, for example with a moving cylinder in a FSI framework, since they would alter its dynamics. In such cases, a sufficient number of particles per cell should be used.
By contrast, when interested in main equilibrium conditions, a $ppc=4$ value appears to provide sufficient solution quality. Numerical experiments show that equilibrium is reached notwithstanding spurious instabilities in the initial transient phase, as shown in temporal averaged pressure in Figure 14. However, as shown in the same figure, a higher resolution allows the accurate representation of small areas, such as the ones in the immediate rear of the cylinder, where quantities, such as the average Mach number, change in a non-negligible way.
Figure 12: Mach 3 flow past a cylinder test case with $N_{C}=10^{4}$, $ppc=4$, pressure field at $t\simeq 1s$.
Figure 13: Mach 3 flow past a cylinder test case, various settings: from the first to the third row, $N_{C}$ changes from $10^{5}$ to $5\cdot 10^{5}$ and $10^{6}$; from the first to the third column, $ppc$ changes from 4, to 9 and to 16.
Figure 14: Mach 3 flow past a cylinder test case, at low resolution (top row, $N_{C}=10^{4}$, $ppc=4$) and high resolution (bottom row, $N_{C}=10^{6}$, $ppc=16$), from left to right: average pressure, pressure standard deviation (represented only to provide a clearer image of the shocks), average Mach number. Values computed as arithmetic average over 85 snapshots corresponding to physical simulation time from 15 seconds to 100 seconds.
4.3 Computational performance
We then focus on the computational performance and scalability of the method, analysing the computational time required by simulations when varying the grid resolution and $ppc$ values. It is worth noting that the implemented parallel kernels can be particle- or grid-based, and, respectively, assign one particle or one grid node to one GPU thread. Moreover, we implemented both mapping kernels (P2G and G2P) as particle-based kernels, since this proved to be the most efficient option.
First, a detailed profiling of each step of algorithm 1 reveals that the G2P and the P2G kernels take up $85\%$ of the parallel part of the code and more than $60\%$ of the whole computational time. This is due both to the relative complexity of those two kernels and to the high number of times they are called per algorithm iteration. We thus focus the scalability analysis on these kernels.
The first finding is that both kernels scale linearly with the total number of particles, in agreement with their implementation (Figures 15 and 16), while pure grid kernels, such as boundary conditions enforcement, scale with the number of nodes. For this reason, we applied a global particles reordering step in the algorithm and studied its effects on performance. Indeed, the ordering is done so that particles which are in the first grid cell after moving are assigned the lowest global indices, followed by those in the second grid cell, and so on moving along the grid. Thus, data which will be accessed at the next time iteration will be ordered, with contiguous particles having contiguous indexes, keeping data locality across the simulation, which is of utmost importance on our target architecture. The reordering turns out to have a significant impact on the G2P kernels, resulting in a further 4x speed-up factor, while preserving the linear scaling (Figure 15).
By contrast, the reordering appears not to have an appreciable effect on the P2G kernels, with slowdowns and accelerations evident at different particle values (Figure 16). The reason can be found in the main implementation difference among the two kernels. P2G introduces a data race, that is currently dealt with using CUDA’s atomicAdd, and its cost becomes more apparent with re-ordering. Indeed, even though, to the best of the authors’ knowledge, NVIDIA’s A100 scheduler policy is not accessible, numerical examples appear to suggest that the scheduler behaves more in a round-robin fashion rather than in a block partitioning way. So, a particle is assigned to a GPU thread, until there are some free, then a second particle is assigned, and so on. In this framework, randomisation of the global order of particles is somewhat beneficial, since it lowers the probability that particles that are in the same cells are processed simultaneously. After reordering, particles close in the domain are also close in memory, enhancing probability of a memory collision, i.e. of simultaneous access to the same data. Since in our case such access would be in read-and-write mode, it activates the atomic operation and degrades performance.
This also explains why reordering appears to not being beneficial for P2G kernels. Since a higher number of ordered particles per cell increases the chance of data races, the average number of particles per cell becomes an important parameter. Indeed, grouping simulations results per $ppc$ value, we recover linearity, and can also confirm the trend of worsening performance at larger numbers of ordered particles in a cell. Indeed, comparing the linear fits in Figure 17, we can see that in the $ppc=4$ and $ppc=9$ cases execution is faster than in the global fit, and in the $ppc=16$ case it is slower.
noindent As for overall computational time, we note that its value largely depends on the requested resolution. On a NVIDIA A100 GPU, to simulate one second of gas flow around the cylinder with $4\cdot 10^{4}$ particles requires about one second of computational time on the device, with $16\cdot 10^{6}$ particles it requires about $3$ minutes. This is reasonably fast, especially considering that the latter resolution is rather excessive for the 2D cylinder test case. For comparison, 3D snow avalanches were recently simulated via MPM with $10$ to $30$ million particles, see [25].
From a previous study [2], we expect that even better performance is achievable. However, with this work we are mostly interested in investigating the potential of MPM when simulating compressible fluids on various GPU accelerated architecture, we thus prioritise portability over efficiency, to be able to compare exactly the same code on different architecture. Portability has been reached also by means of a higher-level approach with respect to low-level CUDA-C, which introduces some level of indirection, with libraries as Thrust that rely on managed memory.
In the next paragraph, we exploit the flexibility granted from this approach to test the code on different hardware architectures, both on CPUs and GPUs.
Figure 15: Profiling of main computational kernels: G2P kernels execution time is compared for the cylinder test case, varying $N_{C}$ and $ppc$, considering also re-ordering effects.
Figure 16: Profiling of main computational kernels: P2G kernels execution time is compared for the cylinder test case, varying $N_{C}$ and $ppc$, considering also re-ordering effects.
Figure 17: Profiling of P2G kernels: effect of average particles per cell number on version with re-ordering.
4.4 Portability
We test the application of the developed implementation on different hardware architectures and compare the performance, addressing the portability of the MPM code. Specifically, we execute the cylinder test case at several resolutions, on different Intel and AMD CPUs and NVIDIA and AMD GPUs, with the same setting described in the previous sections, without enforcing particles re-ordering.
For the first test, we vary the number of cells and of particles per cell and study how this affects the main computational kernels on an AMD CPU, employing $242x$ 24 cores in parallel (Figure 18).
While portability is automatically granted, CPU performance turns out to be worse than single-GPU performance, even though $242x$ 24 CPU cores are employed in parallel. This is consistent with expectations that MPM is better suited to GPU architectures.
As for the computational kernels:
  • P2G kernels scale almost linearly with the number of total particles, in agreement with the beneficial effect of randomisation in reducing the actual execution of atomic operations;
  • G2P kernels maintain quasi-linear scaling only if the average particles-per-cell number is kept fixed;
  • the execution time required by all the other kernels considered together is considerably smaller than the time required by P2G and G2P ones. This confirms the computational relevance of P2G and G2P also on different hardware.
The main difference with respect to the GPU test lies in the comparison between P2G and G2P performance, with the former significantly slower than the latter on the CPU. This is most probably due to our choice of writing an ad-hoc atomicAdd function, based on std::atomic_ref, with the same signature of the more performing CUDA’s atomicAdd, to enhance portability and guarantee parallelism. An algorithm to directly avoid the data-race present in P2G-like kernels, and, thus, to provide a natively portable, parallel and atomic operation free implementation, is proposed in the next section.
Figure 18: Profiling of main computational kernels: total G2P and P2G kernels execution time is compared for the cylinder test case, varying $N_{C}$ and $ppc$, without re-ordering, on $242x$ 24 CPU cores (AMD 7402 CPU).
For the second test, we focus on the heavier test case, the cylinder with $16$ million particles, and simulate one physical second on different architectures, comparing the total wall clock time, referred to the time required by the device, being either a CPU or a GPU; see Figure 19. It is interesting to note that the chosen implementation approach allows the porting of the code to AMD GPUs without any modification.
Figure 19: Comparison of performance on different architectures: overall computational time to simulate 1 physical second of the cylinder test case with $N_{C}=10^{6}$, $ppc=16$.
Lastly, we focus on the comparison between the two GPU architectures.
We can notice that at low computational load, performance is basically equivalent, especially if the higher degree of code optimisation allowed by the NVIDIA compilers is taken into account, even though the AMD GPU seems to be more sensible to the kernels with atomicAdd; see Figure 20.
However, at heavier loads, despite the comparable specs, the AMD MI210 starts to display longer wallclock times than the NVIDIA A100, see Figure 21 for a comparison of the overall computational time on the device. This difference might be due to different aspects, such as having developed the code, and taken performance based design choices, only on NVIDIA A100 GPUs, and having ported it to different hardware afterwards. However, GPU performance is overall competitive with respect to CPU. For example, the heavier test case is 7 to 8 times slower on MI210 than on A100, but anyway 3 to 4 times faster than parallel execution via Intel OneAPI-TBB on Intel Xeon with 2x24 cores.
Most importantly, these results enable us to confirm the portability of the approach.
Figure 20: Comparison of performance on different GPUs: analysis of the average computational time for a low resolution test case $N_{C}=10^{4}$, $ppc=4$, corresponding to about half computational second to simulate 1 physical second.
Figure 21: Comparison of performance on different GPUs: overall computational time required for $Nt=10^{5}$ time loop iterations, varying resolution from $N_{C}=10^{4}$, $ppc=4$ to $N_{C}=10^{6}$, $ppc=16$.
5 Discussion and conclusion
This paper has showcased the potential of a massively parallel, portable implementation of the Material Point Method for compressible flows. Insight has been given on algorithm adaptations which will provide further benefits, moreover, the simulation of standard test cases of compressible flow has provided results comparable with the literature and with established software runs with the same polynomial order (first order shape functions - Lagrangian hat functions) and number of degrees of freedom. These results show that MPM is able to simulate compressible gas dynamics, giving qualitatively good results with a prototype developed modifying existing implementations [65, 14], and serve as further step towards the design of a highly efficient MPM-based transonic and supersonic FSI solver.
As known from the literature [64, 48] and as also noted in a previous implementation [2], $\mathbb{Q}1$ MPM suffers from strong numerical oscillation issues, especially in the vicinity of shocks. The issue is tackled in this paper by adding numerical viscosity in the form of additional terms in the pressure field [63]. A more natural and accurate choice could be obtained by using quadratic B-Spline basis functions [55, 19], a code feature which is currently under development [^8] https://github.com/carlodefalco/quadgrid.
Performance tests on NVIDIA A100 GPUs have confirmed that the main computational steps of the MPM algorithm are the Particles-to-Grid and Grid-to-Particles kernels, and that they scale well with the size of the problem. Moreover, adding a reordering step is beneficial to data locality and, consequently, performance, while data races in the P2G kernel can be tackled CUDA’s atomics at acceptable cost, even though alternative exist. Last, this approach has shown to be easily portable to different hardware architectures, as different vendors GPUs and CPUs.
Several future developments are planned in connection with the implementation and modelling choices made in this paper:
  • On the modelling side, more complex test cases could be simulated, eventually leading to a 3D extension and Fluid-Structure Interaction models. In addition, the effect on parallel performance of higher order B-Splines under development will need to be carefully investigated.
    A 3D extension will facilitate a further growth in computational load, but this will not introduce conceptual main changes in the code, also thanks to the high level, portable approach, which does not refer explicitly to specific, small size memories. However, it could justify a more efficient I/O, e.g., moving from JSON to HDF5. Simulations characterised by a higher computational load would also further benefit from a future multi-GPUs approach.
    The upgrade to a higher order MPM via proper shape function, as B-Splines, would not require a significant change in the numerical method, but will introduce some challenges linked to parallelism.
    In particular, stencil growth due to expansion of shape function support to neighbouring grid elements will introduce halo overlapping, leading to an increase in communication costs, as well as to a constraint on the maximum level of achievable parallel granularity.
    Few works on this topic exist, such as [61], and, since they simulate continuum mechanics materials instead of compressible gases, they can benefit from sparse data structures. However, it is already possible to anticipate that shape function regularity in the current MPM formulation could grow efficiently only up to a certain maximum, such as quadratic or cubic B-Spline, and start to have an impact on performance as shape function support expands.
  • For the algorithm implementation, the use of atomic operations to handle data races was chosen for simplicity, also because it showed an acceptable impact on computational time and scalability on NVIDIA A100 GPUs and a simple portability to other architectures (section 4). To reach the full parallelisation capabilities proved by the G2P kernels also for the P2G ones, P2G algorithm adaptations will be necessary. An attractive and and more portable route would involve avoiding atomics, and writing a grid colouring algorithm that can be implemented in an architecture-independent way, relying only on standard libraries functions. Its main steps are:
    1. 1.
      Build grid-to-particle connectivity
      1. (i)
        Sort particle IDs by cell index
      2. (ii)
        Count particles per cell
      3. (iii)
        Perform a cumulative sum of particle counts
    2. 2.
      Assign a colour to each cell, in such a way that neighbouring cells have different colours; then, for each colour:
      1. (a)
        Perform a for_each on cells
      2. (b)
        Call kernels (12) or (17)
    In this way we would iterate on the vector holding the list of particles in each cell, and then inside each cell we would perform only a small amount of serial operations, without any data race.
  • On the High Performance Computing side, avenues of development include the support of multi-GPU and multi-node architectures via MPI, and a grid-particles adaptive refinement through a quad-tree approach [53, 15]. Moreover, currently the code employs Structure-of-Arrays for the memory, since, differently from arrays of arbitrary structures, it provides coalescing, which improves memory efficiency [30]. However, recent developments on GPUs computing literature [61, 22] show that the best performance is achieved by adopting an Array of SoA (AoSoA). When addressing the MPI-based multi-GPU implementation, this possibility could be explored: indeed, the current data structure gives a conceptually straightforward extension towards multi-GPU implementations, where a further level of parallelisation through an MPI-based domain decomposition will make each GPU hold a SoA as the current one, globally giving rise to an AoSoA data structure.
Acknowledgements
Research was sponsored by the Leonardo Labs, Leonardo S.p.A., through the PhD scholarship “Multi-GPUs based PIC / MPM methods for Compressible Flows” funded at Politecnico di Milano and by ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data, and Quantum Computing funded by European Union - NextGenerationEU. The simulations discussed in this work were, in part, performed on the HPC Cluster of the Department of Mathematics of Politecnico di Milano which was funded by MUR grant Dipartimento di Eccellenza 2023-2027. C. de Falco and P. J. Baioni are members of the GNCS group of INdAM. Views and conclusions contained in this document are those of the author and should not be interpreted as representing the official policies, either expressed or implied, of Leonardo Labs, ICSC, MUR or GNCS-INdAM.
Conflict of interest
The authors declare that they have no conflict of interest.